Find the lowest energy of an electron confined in a cubical box of each side 1 ร…?


1
Expert's answer
2021-03-01T12:38:38-0500

"E=\frac{h^2}{8ma^2},"

"E=\frac{(6.63\cdot 10^{-34})^2}{2\cdot 9.11\cdot 10^{-31}\cdot 10^{-20}}\approx24.12\cdot 10^{-18} ~\text{J}=24.12~\text{aJ}."